Tech Staff Meeting Notes February 2, 2016
The top priority is always repair of down stations.
NM problem stations
ET problem stations
Duty Table updated June 2014
Be sure to send email to ceri_seis for any config changes; and no changes on Friday unless its needed to fix an outage.
We need to have a SIS reboot. Tax season is a good time to enlist John to help populate the tables once the templates are complete.
The TDOT proposal to replace the sensors was funded. VLF has the hardware order.
Server room hardening is underway.
Can we move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire?
USGS water division on south campus needs a new home.
2015 Preventive Maintenance84% complete
NM BB/SM & SP: 67 done & 10 to go
NM SM: 9 done & 8 to go
AG BB/SM: 6 done & 0 to go
GPS: 13 done & 0 to go
Misc: 6 done & 2 to go
freewave serial reset installed at 10 of 11 broadbands. GLAT remains.
Short Period Down
Strong Motion Down
CPAR - spikes
TYAR - spikes
WADM.EHE - spikes
PEBM.HN* - dropping out
PEBM.HH* - dropping out
TNMT and MORT are flooded. Canister at TNMT was pulled and is ready to reinstall. MORT was inaccessible during the last visit.
MORT and MIST traces stopped on 1/21. They are digitized at lnxtworm so should be flat not missing. James will check ew configs. While he's there he will reconfigure the waveservers so that they're on separate disks.
RVAR has slumped down the hill and the pipe vault is now bent. Need to install a repeater to get the outer station, MSAR, back online until access improves enough to rebuild RVAR.
All ISIS3 hardware except that slated for PLAR is installed. Have about 26 remaining, mostly MKTA and some NMAD outer stations.
RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). We have a Basalt almost ready to go so we and install new rather than move old.
NHIN strongmo is down. Jim will see if someone From IU can go down for a recon and check for simple fixes (e.g. fuse).
There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch will check with NSN to see if we need to update it.
ASU museum display needs to have stomp meter finished then be reinstalled.
NCCU was restored with a power cycle. John will draft detailed instructions to help the hosts do this in the future.
The win7 upgrades at nodes (mudworm is done). Steve recommended that unless there are compelling reasons otherwise, we should not upgrade the nodes to Win7. It causes several problems and does not run well on the current hardware. This instigated a further discussion on the obsolescence of the NI card and adsend. John will touch base with previous contacts in Seattle and CA so see if better solutions are farther along.
All three NM nodes and all but two (LPAR and HBAR) NM broadbands got new batteries to replace the ARRA vintage previously there. Have the mktaworm batteries in house but need to be installed.
TDOT funding to replace the senses was approved.
The UPS at aztnworm failed. A temporary fix was installed. Need to finish it up.
Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead.
Data taken from the Russian teleseism:
BCRT.EHN EHN flat
CPRT.HNN long period wobble on HNN
VHTN.EHE ringing on EHE
GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.
John's updated todo list .
Improve RCGA AC power and perform maintenance work on shed.
Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site. Can we use a cell modem?
HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established.
Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.
GFM - Need to do a noise study for a closer site, then move it in the spring along with upgrading to S-13.
Replace VHTN E-W and N-S Instruments.
WMTN pit needs to be moved due to erosion.
We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been formally approved by Greg Hill the center coordinator and the Knox County facility supervisor.
No problems at this time.
FCAR AC power upgrade was abandoned. A tree was removed improving sunlight Temporay additional solar was installed and seems to be working. Need to make it permanent.
Have the hardware; need to build up a new portable for Horton.
Needs some clean up.
The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL
The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.
The seisnet wiki is at https://umwiki.memphis.edu/display/CERITech/CERI+Tech You will need your uuid and password.